Assume that the number of rainy days per month in a given city follows a Poisson distribution with a mean value of 2.4 days. What is the probability that it will rain two or more days next month?
Answer #1

Solution :

Given that ,

mean = = 2.4

Using poison probability formula,

P(X = x) = (e- * x ) / x!

P(X >   2) = 1 - [P(X = 0)- P(X = 1)]

= 1 - (e-2.4 * 2.40) / 0! -(e-2.4 * 2.41) / 1!

=1-0.3084

Probability =0.6916